# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

### v2

@@ 5.2.1. Diverse Decision Strategies Across Visualizations
  • Occlusion Challenges. Some participants (e.g., P2, P7, P11) reported difficulties estimating queue lengths due to occlusion in the X-ray conditions, prompting reliance on distance estimates instead of people counts.
@@ 5.2.2. Increased Trust in AI Suggestions with X-Ray
  • Perceptual and Navigation Challenges. Participants found distance estimation (e.g., P2), queue length assessment (e.g., P4, P5), and locating navigation paths (e.g., P12) more effortful in AR. Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on AI-aided decision-making (2021BansalDoesTheWholeExceedItsParts).
@@ 6.1. Perceptual Challenges of AR X-Ray Can Lead to Over-Reliance on AI
Future Studies. In our study, situational data like queue length was visualized using human-shaped glyphs. A natural possibility is to compare visual encoding, for example, glyphs vs. numeric labels, to evaluate how data representation (e.g., legibility, uncertainty, and saliency) influences AI reliance.
Additionally, examining how different fundamental characterizations of the situation (e.g., task difficulty, information ambiguity, time pressure and many others) modulate trust and reliance in AR embedded visualizations would offer deeper insight into AR-AI interaction design.
@@ 6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
6.3. Aligning the Strengths of AR Embedded Visualizations with Spatial Tasks
In our study, while the X-ray did not outperform the Minimap on decision accuracy, it significantly reduced post-trial pointing errors, suggesting improved spatial mapping. This distinction suggest that AR embedded visualizations may be more beneficial in continuous, action-based spatial workflows rather than in isolated decision-making tasks.
In spatial tasks like emergency response, selecting a correct location is only the first step; what follows is action. Misinterpreting a target’s location after selection can be consequential. For instance, a user might correctly choose an exit from a wall-mounted map but walk the wrong way due to reference-frame confusion. AR embedded visualizations mitigate this risk by spatially anchoring information in the environment. Thus, rather than emphasizing the weakness of AR embedded visualizations, our findings illustrate their potential strengths in embodied spatial tasks and highlight when and where AR embedded visualizations are best applied.
@@ 6.3. Aligning the Strengths of AR Embedded Visualizations with Spatial Tasks
Future Studies. In our study, participants selected targets verbally. A natural follow-up would involve embodied selection, for example, walking to or pointing at targets, to examine whether the spatial support of AR embedded visualizations becomes more critical under physical interaction. We hypothesize that in such embodied tasks, participants using 2D maps will incur greater cognitive load in spatial mapping, possibly diminishing their ability to evaluate targets effectively. Future work should explore how task modality (e.g., static vs. embodied) interacts with visualization and AI support in complex environments.
@@ 6.4. AR X-Ray May Lower Cognitive Engagement Compared with 2D Visualizations
In our study, there was no significant difference in response time between the Minimap+NoAI and X-ray+NoAI conditions once initial search time was excluded. However, participants in the X-ray+AI condition made decisions more quickly yet achieved lower overall accuracy than those in the Minimap+AI condition. This pattern aligns with reliance on heuristic judgments, often described as Type 1 thinking (Chaiklin2012ThinkingFA; JSBTDualProcessTheory2013). Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR embedded visualizations, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
@@ 7. Conclusion
In this paper, we investigated the impact of AR embedded visualizations on human-AI reliance in spatial decision-making under time constraints by comparing X-ray with Minimap. Our user study (N = 32) revealed that the X-ray visualization tends to evoke over-reliance on AI during spatial decision-making. This effect may be linked to perceptual challenges (e.g., occlusion and distance estimation) and AR-specific biases, such as visual proximity illusions and the heightened sense of realism and immersion induced by high-fidelity visuals. Nonetheless, the X-ray visualization exhibits unique strengths by enhancing spatial mapping, as evidenced by significantly reduced pointing errors. We hope that our findings and design implications will guide future research and the development of solutions that better leverage AR’s advantages to improve human-AI collaboration in spatial contexts.

### v3

@@ 5.2.1. Diverse Decision Strategies Across Visualizations
  • Occlusion Challenges. Some participants such as P2, P7 and P11 reported difficulties estimating queue lengths due to occlusion in the X-ray conditions, prompting reliance on distance estimates instead of people counts.
@@ 5.2.2. Increased Trust in AI Suggestions with X-Ray
  • Perceptual and Navigation Challenges. Participants found distance estimation (P2), queue length assessment (P4, P5), and locating navigation paths (P12) more effortful in AR. Thus, AI suggestions offered a useful cognitive shortcut, aligning with findings from prior research on AI-aided decision-making (2021BansalDoesTheWholeExceedItsParts).
@@ 6.1. Perceptual Challenges of AR X-Ray Can Lead to Over-Reliance on AI
Future Studies. In our study, situational data like queue length was visualized using human-shaped glyphs. A natural possibility is to compare visual encoding, for example, glyphs vs. numeric labels, to evaluate how data representation like legibility, uncertainty and saliency influences AI reliance.
Additionally, examining how different fundamental characterizations of the situation, such as task difficulty, information ambiguity and time pressure, modulate trust and reliance in AR embedded visualizations, would offer deeper insight into AR-AI interaction design.
@@ 6.2. Cognitive Biases Triggered by AR X-Ray Can Induce Inappropriate AI Reliance
6.3. Aligning the Strengths of AR X-Ray with Spatial Tasks
In our study, while the X-ray did not outperform the Minimap on decision accuracy, it significantly reduced post-trial pointing errors, suggesting improved spatial mapping. This distinction suggest that AR X-Ray may be more beneficial in continuous, action-based spatial workflows rather than in isolated decision-making tasks.
In spatial tasks like emergency response, selecting a correct location is only the first step; what follows is action. Misinterpreting a target’s location after selection can be consequential. For instance, a user might correctly choose an exit from a wall-mounted map but walk the wrong way due to reference-frame confusion. AR X-Ray mitigates this risk by spatially anchoring information in the environment. Thus, rather than emphasizing the weakness of AR X-Ray, our findings illustrate their potential strengths in embodied spatial tasks and highlight when and where AR embedded visualizations are best applied.
@@ 6.3. Aligning the Strengths of AR X-Ray with Spatial Tasks
Future Studies. In our study, participants selected targets verbally. A natural follow-up would involve embodied selection, for example, walking to or pointing at targets, to examine whether the spatial support of AR embedded visualizations becomes more critical under physical interaction. We hypothesize that in such embodied tasks, participants using 2D maps will incur greater cognitive load in spatial mapping, possibly diminishing their ability to evaluate targets effectively. Future work should explore how task modality, such as static vs. embodied, interacts with visualization and AI support in complex environments.
@@ 6.4. AR X-Ray May Lower Cognitive Engagement Compared with 2D Visualizations
In our study, there was no significant difference in response time between the Minimap+NoAI and X-ray+NoAI conditions once initial search time was excluded. However, participants in the X-ray+AI condition made decisions more quickly yet achieved lower overall accuracy than those in the Minimap+AI condition. This pattern aligns with reliance on heuristic judgments, often described as Type 1 thinking (Chaiklin2012ThinkingFA; JSBTDualProcessTheory2013). Together with the high self-reported confidence, these results suggest a lack of deeper cognitive engagement: once AI assistance was available, participants quickly felt confident enough in their intuitive judgments, rather than engaging in more analytical (Type 2) thinking. These findings imply that AR X-Ray, by offering a highly immersive or embodied experience, may inadvertently reduce users’ motivation to scrutinize information.
@@ 7. Conclusion
In this paper, we investigated the impact of AR embedded visualizations on human-AI reliance in spatial decision-making under time constraints by comparing X-ray with Minimap. Our user study (N = 32) revealed that the X-ray visualization tends to evoke over-reliance on AI during spatial decision-making. This effect may be linked to perceptual challenges like occlusion and distance estimation, and AR-specific biases such as visual proximity illusions and the heightened sense of realism and immersion induced by high-fidelity visuals. Nonetheless, the X-ray visualization exhibits unique strengths by enhancing spatial mapping, as evidenced by significantly reduced pointing errors. We hope that our findings and design implications will guide future research and the development of solutions to better leverage AR’s advantages to improve human-AI collaboration in spatial contexts.